### v1

Abstract
We explore the enhancement of Human-in-the-Loop video annotation by integrating automatic capabilities to ease the task for annotators and assess their performance. The research delves into the practical implications of the annotation processes, the integration of AI components, and the evaluation of its outcomes. We analyze their impact on efficiency, accuracy, and overall annotation quality.
Focusing on the Human-in-the-Loop for video annotation tasks, we implemented a single-iteration scheme using Label Studio and AI-powered zero-shot pre-annotations. Using this framework, we designed a test based on the annotation of the UCF-Crime dataset to discriminate between normal and abnormal activities in video footage.
Our results evidence how automatic AI-based pre-annotation can streamline the video annotation workflow, empowering human annotators and optimizing the overall pipeline. Using the pre-annotated data, we observed a 35% reduction in the annotation time for 70% of the annotators with similar quality annotations, compared to the traditional manual annotation task. Results are consistent with asset duration and complexity. We also observed that while annotators rapidly learned to use the tool, the produced annotations are more coherent among annotators and better match the natural clustering of the video frames.  [^1] The collected data is accessible on https://github.com/jgnav/AI-Boosted-Video-Annotation/.
1 Introduction
Video annotation has been the cornerstone of data-driven learning, particularly in scenarios demanding highly reliable annotated data. Alleviating the burden for annotators, expediting the process, and improving the annotations are recurrent goals to which we want to contribute.
By incorporating Artificial Intelligence (AI) tools into the annotation process within the Human-in-the-Loop (HITL) paradigm, the aim is to reduce the number of frames to annotate and make each human input more valuable, reducing the work time and minimizing the probability of an error or inconsistency. We are leveraging human intelligence to handle the nuances and complex decisions, which are the weak points of fully automated annotation, and harnessing the machines’ ability to process large amounts of data.
We devised a comprehensive annotation test to assess pre-annotations’ impact on the video annotation flow. We aim to explore the advantages and disadvantages of this approach and its impact on final data outcomes. We have decided to conduct the test using real-world videos primarily captured from surveillance cameras. Surveillance videos are characterized by their typically low resolution and poor quality, recorded from a stationary point amidst varying lighting and environmental conditions. This choice is because it presents a challenging use case for pre-annotating automatically with AI models and is a widely valued scenario in today’s context.
Additionally, we chose a group of annotators for data labeling instead of relying on a limited set of “experts”. This approach leverages the collective intelligence of a diverse pool of individuals for tasks such as labeling, classification, and data annotation. Nonetheless, it would be interesting to evaluate the results of a similar test on professional annotators to compute performance figures and evaluate their proficiency level. For the time being, this was out of the scope of our investigation.
Crowdsourcing annotations for discrete labeling has become a major trend these days. The task involves collecting multiple judgments from annotators, which are then aggregated to establish a single “ground truth” published in the dataset [6]. However, it is crucial to recognize that disagreements among annotators may contain valuable nuances worth exploring. Our test follow a crowdsourcing methodology, implementing stringent quality control measures to ensure the generation of high-quality annotations. These measures include mandatory training on the annotation tool before the test, removing outliers (annotations that deviate logically from the expected average), and assigning each video for annotation with a 6-vote overlap. This ensures that each task receives input from six different annotators, promoting diversity in perspectives.
Therefore, our primary objective is to design a hypothetical annotation test that affords annotators significant freedom, allowing for the consideration of diverse criteria. Subsequently, we conducted a study on and analyzed the impact of our pre-annotations on annotators’ criteria and efficiency and evaluated the consistency of the results obtained.
2 Related work
The literature describes several different annotation tasks and platforms. Reviewing previous initiatives incorporating AI-based pre-annotations to streamline the annotation process, we observe the inherent complexity and diversity, adding a layer of intricacy to the evaluation. However, let us first clarify the concept of data annotation: Data annotation involves the labeling or tagging data with specific labels or tags, rendering it more comprehensible for machines. This crucial activity is pivotal in developing supervised Machine Learning (ML) algorithms and models and validating their unsupervised and self-supervised counterparts [26].
These annotations vary significantly in terms of the task at hand and the tools used, and these aspects must be considered in the annotation process.
2.1 Common annotation tasks
Video assets represent rich sources of information, prevalent across a diverse range of applications and tasks associated with information management. The extensive variety of video annotation tasks contributes to a broad spectrum of annotation tasks.
The granularity required for video annotation gives rise to many annotation cases. The annotation may encompass the entire video [21, 27], temporal segments (shots) [27, 1], frames (temporal segments with zero duration) [9, 32], regions within frames [20, 19], or even moving regions—referring to a region tracked across a sequence of frames [4].
Another important use case relates to the Video Moment Retrieval (VMR). VMR retrieves a segment described by a given natural language query from an untrimmed video. This task, alternatively known as natural language video localization and video temporal grounding, is a fundamental challenge in comprehending computer vision and visual information retrieval. VMR finds wide applications in scenarios such as video browsing websites and semantics-based video search engines [30, 3, 2]. Interestingly, the annotations introduced for each use case vary significantly, and therefore, we must consider the task while looking for AI tools that may boost the annotation process.
2.2 Tools for video annotation
We have reviewed the characteristics of the primary video annotation platforms. CVAT, an open-source annotation platform developed by Intel, offers a comprehensive set of annotation tools, including object detection, image segmentation, and video annotation. CVAT supports multiple annotation formats and provides APIs for easy integration with other computer vision applications and workflows. Being a web-based tool, CVAT can be downloaded and installed on custom local or cloud servers, making it customizable and extensible. It also includes tools for medical image conversion and annotation [16].
ANU CVML VIDeo Annotation Tool (VIDAT) is a high-quality, simple, and efficient open-source video annotation tool developed by ANU Computer Vision Machine Learning. It is user-friendly, supporting multiple annotation types, including temporal segments, object bounding boxes, semantic and instance regions, tracklets, and human pose (skeleton). VIDAT runs in a browser without external libraries or the need for server-side processing. However, it is easy to integrate with a back-end “Human-in-the-Loop” processing. It ensures data security as the data does not need to leave the local machine, and it is open-source [14].
The Video and Image Annotation (VIA) is a lightweight, standalone, and offline software package that operates in a web browser without requiring installation or setup. Users can annotate at various levels of granularity, from the entire video down to specific segments, moving regions, individual frames, and even static regions within a frame. VIA also supports collaborative annotation of a large dataset by a group of human annotators. The BSD open-source license allows its use in academic projects or commercial applications [7].
The IBM VideoAnnEx annotation tool focuses on video annotation utilizing MPEG-7 metadata. Despite the tool no longer being actively maintained, VideoAnnEx remains accessible and serves as an illustrative example of content annotation aligned with the MPEG-7 initiative. While its maintenance status may pose considerations for long-term use, the tool offers valuable insights and a historical perspective on content annotation methodologies within the MPEG-7 framework [15].
Finally, despite the features described, our choice for the tests was Label Studio for several compelling reasons. Firstly, Label Studio has proven to be a reliable platform with a track record of annotating high-quality data. It is free and open-source and allows extensive customization options for the annotation interface. Moreover, it provides essential functionalities for our test, such as frame classification over a temporal bar and the automatic, precise measurement of annotation time per video [24].
2.3 AI-boosted annotation
Video annotation constitutes a time-consuming and labor-intensive endeavor, prompting the exploration of various approaches to enhance efficiency through integrating the Human-in-the-Loop paradigm. The primary goal is to reduce the number of frames requiring annotation, thus maximizing the value of each human input in the process. The expansive landscape of video annotation applications has given rise to diverse methodologies tailored to specific use cases.
In [13], a unified framework for generic video annotation with bounding boxes is introduced. This model combines interpolating and extrapolating capabilities, tracking target appearance in multiple annotated frames. Additionally, it proposes a guiding mechanism that sequentially generates suggestions for the next frame based on previous annotations.
The work presented in [3] focuses on video moment retrieval, aiming to identify the start and end timestamps of a moment in a video based on a natural language query. The introduced paradigm, “glance annotation”, requires only the timestamp of a single random frame, referred to as a “glance”, within the temporal boundary of the fully supervised counterpart. ViGA, utilizing contrastive learning, divides the input video into clips and contrasts them with queries, assigning glance-guided Gaussian distributed weights to all clips.
In [5], a data analysis framework is proposed to facilitate the interactive annotation of racket sports videos with the support of computer vision algorithms. This framework leverages machine learning models in computer vision to help users extract essential events from videos and provides interactive tools for data annotation.
A lightweight approach named Video2Action is presented in [8]. In video creation, where annotating actions is a common requirement, Video2Action automatically generates action scenes and predicts action locations in the video using image processing and deep-learning methods.
Other authors have developed their AI-powered tools to ease video annotation. One of these is the pre-annotation tool developed by the authors of the present contribution using contrastive learning and cross-modal representations (to be published).
3 Annotation Assessment
The annotation process focused on classifying frames from surveillance camera videos containing anomalous events. This task requires annotators to categorize and group the contents of each video. It is important to acknowledge that the criteria for classification may vary among annotators, and there may not be a single valid criterion, even though we defined a single task.
Annotators interacted with a video interface that featured the video itself and a temporal bar, facilitating the convenient classification of frames into five distinct classes. The interface also includes checkboxes for binary selection on each class, allowing annotators to indicate whether the video segment corresponds to an abnormal event (see Figure 1).
Figure 1: Label Studio interface [24] as adapted for the performed experiences and test.
Our annotation pipeline generates pre-annotations, providing comprehensive annotations for each video. Additionally, keyframes representing each class serve as visual references during the annotation process. Therefore, there are two types of annotations: without pre-annotation or unassisted, and with pre-annotation. In the second case, the annotator is presented with the fully pre-annotated video, and their task is to review the pre-annotation and modify it as they deem fit according to their criteria. We generated these pre-annotations using multi-modal AI models incorporating both vision and text. This work is still pending publication.
Our objective is to study the differences in annotations with and without pre-annotation and to identify any potential biases introduced by pre-annotations. Another crucial aspect under consideration is annotation time; our pipeline aims to improve the annotation process’s efficiency significantly.
For a more in-depth description of the annotation process, our CrowdWorkSheets responses [6] have been included in Appendix 0.A, a framework widely employed in similar works [10].
3.1 Experiment setup
We selected Label Studio as our annotation platform for this study [24]. To carry out the annotation process, we engaged the participation of 18 voluntary annotators, all affiliated with a university. Each annotator was assigned ten different videos, with a unique combination of five videos containing pre-annotations and five without pre-annotations, presented in varying orders. Half of the annotators began with pre-annotated videos, while the other half started with videos without pre-annotations. The approximate time required per annotator was 30 minutes, excluding training. In total, 30 videos underwent annotation, each receiving six annotations from six different annotators (6-vote overlap) - three with pre-annotations and three without.
The video selection process involved a random sampling from all predefined annotation classes. Two filters were applied during the selection: the first filter was based on video duration, discarding those lasting over 40 seconds due to time constraints for annotators. We aimed for a diverse set of videos with shorter durations. The second filter excluded videos with explicit content, determined by the consensus of three researchers.
The entire annotation process was overseen by Sigma Cognition S.L. [11], an entity with extensive experience in annotation tasks. Their supervision ensured the quality and consistency of the annotations throughout the project.
3.2 Dataset of Real-world Videos
The dataset chosen for this task is UCF-Crime, which is recognized as one of the most comprehensive and diverse collections of surveillance videos. This dataset has been extensively utilized in research on this specific type of video domain [28, 12, 31].
The dataset comprises a total of 128 hours of videos, consisting of 1900 long untrimmed real-world surveillance videos. It covers 13 realistic anomalies, including fights, traffic accidents, burglaries, robberies, and more. These anomalies are significant due to their impact on public safety [22].
4 Results
An extensive analysis of the annotation test results has been conducted and divided into three main blocks. Figure 2 illustrates the framework of all the result evaluation processes carried out, including the inputs and outputs of each. Two primary evaluation metrics are extracted from the annotation process: frame classification and the annotation time per video for each annotator. Concerning the temporal analysis, the objective has been to measure if there has been a substantial gain in annotation time.
Within the quality evaluation block, an assessment was made to determine whether there was a noticeable change in the quality of annotations, examining their consistency and the potential introduction of biases. Finally, in the evaluation block utilizing embeddings representing the semantic information of the frames, the classification quality is assessed about the video’s semantic content.
Pre-anno-tationsAnnotationProcessTimestampsTemporalEvaluationTemporalAnalysisQualityAnalysisVisualSimilarityAggregation&HomogeneitySemanticLatent SpaceOperationsLabeled ClassesFrames EmbeddingsCLIPImageEncoderAI-BasedPre-annotationPipelineVideos
Figure 2: Diagram of the process followed to evaluate the results of the annotation test and to run the experiences. Videos (to the left) are automatically pre-annotated. These pre-annotations help the annotation process, while input frames are processed using the CLIP image transformer [18]. Annotation results (timestamps, labeled classes) and frame embeddings are used to assess annotation results.
4.1 Temporal analysis of annotations
The most crucial metric in this test is the temporal one, as the primary goal is to optimize the annotation time. To measure this, the annotation platform Label Studio accurately tracks the time elapsed from when the annotator opens the task until they click the submit button.
Given the number of annotations obtained, it is crucial to consider specific factors that may influence the temporal metric, especially concerning the presence of human factors and video-related elements in the data. Nonetheless, our focus is on studying discernible trends. More issues that were unrelated to the test, such as technical problems, annotator distractions, or those arising from inexperience, have been addressed separately by identifying outliers and assessing whether it is worthwhile to eliminate them from the annotations.
Figure 3 illustrates the distributions of all annotation data. The kernels of these distributions exhibit a behaviour similar a Gamma functions. Table 1 provides a comprehensive overview of the statistical analysis results. Parametric tests were carried out between the two annotation sets to evaluate their independence, and all tests affirmed their independence. On average, the annotation time is lower by 23.11% for cases where pre-annotation is applied.
Figure 3: Distribution of annotation times with and without pre-annotations. Bin steps of 20 seconds were employed. The kernel density estimation plot illustrates the trend for both distributions.
Table 1: Analysis of the independence between two sets through parametric statistical testing. The outcomes of the statistical test for independence are detailed in the upper section. The lower section provides a statistical overview of the two distributions.
Parametric test Statistic $p$-value Satisfies independence assumption ($p$ << 0.005)?
$t$-test -3.6391 0.0004 True
One-way ANOVA 13.2428 0.0004 True
Wilcoxon signed-rank 853.0000 0.0002 True
Annotation time
Type of annotation #Annotations #Videos Mean (s) STD (s)
With pre-annotation 90 30 146.42 79.28
Without pre-annotation 90 30 190.44 72.63
Comparing annotation times per video is less meaningful than per annotator since human factors exert the most significant influence. Figure 4 illustrates that 72% of annotators took less time to complete their annotations with pre-annotation, with a median improvement of 35%. While acknowledging potential influences from human factors and video content, efforts have been made to isolate annotation times as much as possible. Figure 5 explores the relationship between annotation times, the number of classes per video, and their duration, revealing no direct correlation.
Annotator feedback suggests that the complexity of video annotation plays a significant role in annotation time, involving separable actions, content comprehension, and video quality. Although challenging to quantify, annotators have concurred that pre-annotation significantly reduces decision-making time regarding class distribution. Multiple cases of “perfect” pre-annotation were reported, where annotators only needed to review and submit the pre-annotated results, resulting in maximum time savings. Hence, analyzing raw data trends, most annotators experienced reduced annotation time, indicating tangible efficiency gains.
Figure 4: Comparison of annotation times per anonymized user. Each point corresponds to the annotation times of an annotator, and the median of all annotation times with and without pre-annotation has been taken. Figure 4(a) contains the total annotation times per video, while in figure 4(b), these times are normalized by the duration of each video. Points above the green area indicate that the median annotation time with pre-annotation is shorter than without, whereas points above the red area suggest the opposite.
Figure 5: Figure 5(a) displays the median annotation time for all videos grouped by the number of non-unique classes. Figure 5(b) is the median annotation time per video duration, with intervals set at 5 seconds. In both cases, the data is separated by annotation type.
Lastly, a phenomenon extracted from the data is the impact of training. According to annotator feedback, pre-annotation utilization improved with experience in annotation. We removed the first four annotations from each annotator to observe these effects and compared the results with those obtained in Figure 4. Figure 6 highlights the vectors depicting the differences. Most annotators benefited more from pre-annotations in their last six out of ten.
Figure 6: Comparison of results obtained in Figure 4, excluding the first four annotations. Each red arrow represents the change for each annotator. The farther from the boundary between the zones, the greater the increase in annotation time difference with and without pre-annotation.
4.2 Quality Analysis of Annotations
The quality analysis has two components: we attempt to estimate the ground truth through our aggregation methodology, and the second involves comparing the annotations among themselves. The intrinsic freedom of the test does not allow for a broad consensus on what is considered a “correct annotation”.
When aggregating such annotations, it is crucial to consider that labels are independent of absolute values, and a permutation of labels should not alter the result. Only the clustering of frames is valued. This consideration is essential both during aggregation and when selecting evaluation metrics. The methodology for obtaining our ground truth is as follows:
  • Extract all annotations of the same type from each frame.
  • Create a tuple with the annotations and hash it. Using hashing in this context helps create a unique identifier for each different group of frames without relying on the specific values of the labels. This approach allows us to handle cases with an infinite number of classes.
  • The new labels are the hashes of each group, considered as the ground truth. This way, we preserve only the parts where annotations coincide, even in the presence of noisy annotations.
  • Calculate metrics tailored to our use case with the ground truth and each annotation, averaging all combinations.
This approach helps address the noise issue and enables the study of patterns of similarity in clustering. Considering the hashes of frame annotations as the ground truth results in clusters where all annotations have deemed these frames to be grouped similarly. Therefore, we can consider these clusters as “strong clusters”, while the rest of the information is discarded.
The next step involves selecting metrics that measure only the shared information between the two classifications, disregarding the absolute value of the label, permutation of classes, and the number of classes. The selected metric is Adjusted Mutual Info (Adjusted, Normalized and V-measure) defined by the equation:
$\text{Mutual Information}=\sum_{i=1}^{\mathbf{|U|}}\sum_{j=1}^{\mathbf{|V|}}\frac{\mathbf{|U_{i}\cap V_{j}|}}{\mathbf{N}}\log\left(\frac{\mathbf{N|U_{i}\cap V_{j}|}}{\mathbf{|U_{i}||V_{j}|}}\right).$ (1)
Here, $\mathbf{|U_{i}|}$ represents the number of samples in cluster $\mathbf{U_{i}}$, $\mathbf{|V_{j}|}$ represents the number of samples in cluster $\mathbf{V_{j}}$, and $\mathbf{N}$ is the total number of samples. Table 2 presents the results, indicating no significant difference in the classification quality.
Table 2: Statistics from the test comparing annotations of different types with the aggregated ground truth.
Type of Annotation Adjusted Mutual Information Normalized Mutual Info V-measure
With pre-annotation 0.64 0.64 0.64
Without pre-annotation 0.63 0.64 0.64
Another aspect to explore is whether a discernible trend toward annotation homogenization exists when utilizing pre-annotation. Since a lot of freedom was given in annotations precisely to deduce the effect of pre-annotation, Table 3 contains the results of our homogenization test. Metrics were calculated with all combinations of pairs of annotations of the same type for each video, and the mean of all these combinations is represented. As observed in the table, the similarity between annotations with pre-annotation is higher, indicating greater homogeneity. This capability empowers designers to steer outcomes by leveraging pre-annotations to attain specific annotation nuances or uphold quality standards, thereby mitigating annotator fatigue.
Table 3: Statistics from the test on the homogeneity of annotations of different types.
Type of Annotation Adjusted Mutual Information Normalized Mutual Info V-measure
With pre-annotation 0.67 0.67 0.67
Without pre-annotation 0.62 0.62 0.62
4.3 Visual Similarity: Evaluating Results with CLIP
The CLIP [18] similarity (denoted as CLIP-SIM) metric has been widely employed in various studies [33, 23, 29, 17] to assess image generation models primarily. However, we aim to adapt this metric for video evaluation, following a methodology similar to FVD [25]. After encoding frames with the “Vit-32” image encoder from CLIP, we calculated the Silhouette Score using the cosine distance. This metric, in conjunction with embeddings and a distance function between them, aims to estimate the quality of clustering based on the semantic content of frames.
As illustrated in Figure 7, we observed a higher Silhouette Score in 25 out of 30 pre-annotated videos. The mean Silhouette Score with pre-annotation was 0.41, whereas without pre-annotation, it was 0.28, indicating a significant difference of 52.10%.
These results are expected since pre-annotation involves starting with a frame classification that maximizes the Silhouette Score. Considering this metric as an objective measure of classification quality, which respects the semantic content of the video frames, it can be inferred that pre-annotation enhances the annotation’s overall quality.
Figure 7: Comparison of the median Silhouette Score for all annotations of the same type within each video. Each point represents a video, totaling 30 points. Points in the green area indicate that the median Silhouette Score of pre-annotated annotations is higher than without annotation, and vice versa for the red area.
5 Conclusions
We have successfully implemented the Human-in-the-Loop paradigm for video annotation using automatic pre-annotations and conducted a full test to evaluate the annotation process, comparing its results with and without using these AI-boosted pre-annotations.
We observed that 70% of our novel annotators achieved a 35% reduction in their annotation time using the proposed HITL annotation scheme while delivering similar annotations.
We observed that the annotations differed, yet they remained equally valid. This divergence may relate to the freedom provided in the test, and even when analyzing the semantic content using the CLIP vision model, an improvement in classification concerning semantic content was noted. This suggests that introducing a bias that can be advantageous or detrimental depending on the test design. Moreover, it can be considered an additional factor in shaping the output of annotations, such as determining the level of detail or the type of actions to be classified.
Another valuable and qualitatively unmeasurable aspect throughout the process was the feedback from the annotators. They expressed that the pre-annotations reduced annotation fatigue, facilitating easier annotation with the tool. Such insights gain significance when extrapolated to routine annotation projects that entail hundreds of hours and involve a larger number of annotators.
Finally, the methods developed to conduct these experiments and the technology used in the pre-annotation have been devised to ease the annotation process. Nonetheless, the AI boost in pre-annotation may impact the workforce currently performing annotation duties, affecting the whole sector. First, the proposed technology may reduce dependence on human labor, ultimately relegating annotators to the most complex and challenging cases. Second, these technologies raise multiple concerns regarding the digital rights of the data and the annotations and how to preserve ownership and control it. From a technological standpoint, we understand that digital transformation will also affect the data annotation market. Evaluating the capabilities covered in this contribution sets the path for a deeper and broader discussion. More importantly, it will provide evidence of the pivotal role of the AI-powered HITL in transforming the data industries.
Acknowledgments
We appreciate the volunteer annotators for their dedication and time, as well as the Label Studio developers for providing an open-source annotation platform for this project.
Research framed under HADA project “Research on advanced data annotation tools" (2021/C005/00146323), funded by EU Next Generation (through Ministry of Economy and Digital Transformation & RED.ES).
Disclosure of Interests. Juan Gutiérrez received research funding at UPM from a project funded by Sigma Cognition.
Appendix 0.A Experimental evaluation CrowdWorkSheets
[As supplemental material, we include further details of the video annotation test that was conducted. This material should be published along with the manuscript and the results obtained.]
To further describe our annotation process, we include our responses to the CrowdWorkSheets questionnaire [6]. CrowdWorkSheets is a framework for transparent and ethical dataset annotation that emphasizes the annotation process’s documentation, considering the annotators’ identities and experiences. The framework provides guidance on ethical considerations, annotator selection, and documentation of the annotation process, aiming to address unresolved ethical issues in crowdsourcing. More importantly, it allows us to share a complete summary of the annotation process conducted to ease replicating the test and to highlight the key aspects behind its execution.
0.A.1 Task Formulation
At a high level, what are the subjective aspects of your task?
This annotation test is highly subjective in various aspects. The annotator can decide on the criteria for grouping the frames. This decision has been made intentionally in the test’s design, as it aims to contrast with the classification criteria of our pipeline using CLIP. Determining what constitutes an “anomalous” or “normal” event is quite ambiguous, even for individuals. The most effective way to achieve a valid evaluation is through consensus among multiple annotations.
Another significant point of subjectivity is the delineation between events. Sometimes, the boundary between the end of one event and the beginning of the next, or even whether they can be considered separate events, is very subtle. To minimize the bias of each annotator and select the most “valid” annotation, we have employed a method that involves obtaining three annotations for each type from each video. Consequently, we gathered six annotations for each video, providing greater confidence in the evaluation.
What assumptions do you make about annotators?
Some of the key assumptions we made about our annotators are as follows:
  • Providing the context of the test and explaining the objectives of this work, where disagreements are expected, is intended to enhance each annotator’s criteria to achieve diverse results.
  • We assume that annotators are committed to conducting the test accurately. This assumption is based on their status as trusted individuals within a university department and their academic interest in the test, given their involvement in projects of the same theme.
  • It is assumed that all annotators possess their own personal criteria for evaluating what is considered anomalous, acquired through lived experiences.
How did you choose the specific wording of your task instructions? What steps, if any, were taken to verify the clarity of task instructions and wording for annotators?
To delineate the task without overly restricting the annotator’s freedom of judgment, a somewhat broad definition of anomalous events has been provided: criminal, accidental, or catastrophic events. The instructions have been validated multiple times by different researchers, incorporating modifications.
To help the understanding of any annotator, instructions start from the lowest possible level, using standard, concise, and precise vocabulary. Additionally, the test was deployed in groups, beginning with the most confident annotators, and the annotation guide was updated after each batch.
Any annotator who requested it has received direct support to clarify any doubts that may have arisen during the process. This approach ensures a clear and consistent framework for the annotation task.
What, if any, risks did your task pose for annotators, and were they informed of the risks prior to engagement with the task?
Our task involves annotators reviewing videos that potentially contain violent, explicit, catastrophic, and other types of harmful content. Consequently, the task carries a risk of psychological harm to annotators. Additionally, we have deliberately selected annotators who share cultural similarities with the researchers to ensure that, after filtering the videos, the criteria for identifying images considered too intense for annotators are aligned. Therefore, annotators from significantly different nationalities and cultures from the Spanish context were excluded.
Before commencing the task, each annotator was informed of the video content to be encountered, ensuring they were adequately prepared and provided their consent. Furthermore, the videos underwent an additional review by different researchers, allowing each to exercise discretion in excluding content deemed excessively intense for the annotator.
What are the precise instructions that were provided to annotators?
Annotators will be tasked with annotating crime videos captured by surveillance cameras. The estimated duration of the test is less than 30 minutes, with 2-4 minutes allocated for each video. Prior filtering has been conducted to exclude videos containing overly sensitive imagery.
Each annotator is required to annotate 10 videos, with 5 left without pre-annotation and 5 pre-annotated automatically. In the case of pre-annotated videos, annotators are instructed to review the annotations and make necessary modifications.
Annotation involves grouping frames with similar content by marking them on a temporal bar and assigning a class to each segment. The criteria for considering content as similar include the occurrence of the same action, any event, or a similar setting. Additionally, each segment should be classified as either a “normal event” or an “abnormal event”, the latter being defined when a criminal, accidental, or catastrophic event occurs.
Each annotator has been assigned a profile within the annotation tool, specifying the order in which the assigned videos should be annotated. We provided the annotators with a guide for using the annotation tool interface has been provided, along with tips on navigating the selection bar, using commands, and assigning classes.
It is recommended that annotators practice with a sample video before becoming familiar with the tool. Additionally, annotators are advised to view the entire video before annotation to grasp the context of all events occurring globally in the video.
0.A.2 Selecting Annotators
Are there certain perspectives that should be privileged? If so, how did you seek these perspectives out?
No. N/A
Are there certain perspectives that would be harmful to include? If so, how did you screen these perspectives out?
No. N/A
Were sociodemographic characteristics used to select annotators for your task?
The annotators are all volunteers who chose to participate in the task. They are exclusively Spanish individuals with a minimum academic background of university education. None of the annotators have prior experience in annotation tasks, and they are all affiliated with a university. The annotator pool is well-balanced in terms of gender, and the age range spans from 18 to 50.
If so, please detail the process. If you have any aggregated sociodemographic statistics about your annotator pool, please describe.
Other sociodemographic statistics about our annotator pool were not available.
Do you have reason to believe that sociodemographic characteristics of annotators may have impacted how they annotated the data? Why or why not?
It is possible that, being a somewhat homogeneous group, annotators may share a similar perception of what is considered normal or abnormal in surveillance camera videos.
Consider the intended context of use of the dataset and the individuals and communities that may be impacted by a model trained on this dataset. Are these communities represented in your annotator pool?
The purpose of this test is solely to evaluate our pre-annotation pipeline. Although the potential bias introduced into the data by pre-annotations is studied in 4.2.
0.A.3 Platform and Infrastructure Choices
What annotation platform did you utilize?
We used Label Studio [24] for our annotation needs.
At a high level, what considerations informed your decision to choose this platform?
We opted for Label Studio for several reasons: Firstly, it is a generally reliable platform with a track record in annotating high-quality data. It is free and open-source, providing extensive customization options for the annotation interface. Additionally, it offers essential functionalities for our test, such as frame classification over a temporal bar and automatic, precise measurement of annotation time per video.
Did the chosen platform sufficiently meet the requirements you outlined for annotator pools? Are any aspects not covered?
The selected platform adequately covered all the needs of this experiment.
What, if any, communication channels did your chosen platform offer to facilitate communication with annotators? How did this channel of communication influence the annotation process and/or resulting annotations?
Both direct interaction and an external medium, email, were utilized for communication. Technical queries regarding the tool were addressed directly. In uncertainty related to the annotation process, annotators were informed of the freedom of interpretation, emphasizing that no criterion is more valid than another to avoid influencing annotations.
How much were annotators compensated? Did you consider any particular pay standards, when determining their compensation? If so, please describe.
No compensation was provided to any annotator; all participated voluntarily. While some may have academic interest in this study, no financial incentives were offered.
0.A.4 Dataset Analysis and Evaluation
How do you define the quality of annotations in your context, and how did you assess the quality in the dataset you constructed?
To ensure the precision of annotation time in a hypothetical scenario involving the annotation of large volumes of data, annotators initially practiced using a test video to familiarize themselves with the annotation tool. Outliers were eliminated if they exhibited a significant deviation from the rest of the annotations. Assessing the accuracy of annotations in this test is challenging due to the experimental nature and lack of a ground truth. In such cases, the only viable approach is to annotate the same video multiple times to determine the average consensus among annotations.
Have you conducted any analysis on disagreement patterns? If so, what analyses did you use and what were the major findings? Did you analyze potential sources of disagreement?
An extensive study on the consistency of annotations and, consequently, the patterns of disagreement has been conducted; it is detailed in [citation].
How do the individual annotator responses relate to the final labels released in the dataset?
As mentioned earlier, the release of dataset annotations is not planned at this stage, as we are currently evaluating the annotation process. However, in a real annotation scenario, the classification of frames and keyframes would be made public, employing an agreement method among annotators’ classifications. Therefore, each annotator’s annotations would serve as a vote in the agreement process.